# Continuous Diffusion for Mixed-Type Tabular Data

### v3

@@ Abstract
Score-based generative models (or diffusion models for short) have proven successful for generating text and image data. However, the adaption of this model family to tabular data of mixed-type has fallen short so far. In this paper, we propose CDTD, a Continuous Diffusion model for mixed-type Tabular Data. Specifically, we combine score matching and score interpolation to ensure a common continuous noise distribution for both continuous and categorical features alike. We counteract the high heterogeneity inherent to data of mixed-type with distinct, adaptive noise schedules per feature or per data type. The learnable noise schedules ensure optimally allocated model capacity and balanced generative capability. We homogenize the data types further with model-specific loss calibration and initialization schemes tailored to mixed-type tabular data. Our experimental results show that CDTD consistently outperforms state-of-the-art benchmark models, captures feature correlations exceptionally well, and that heterogeneity in the noise schedule design boosts the sample quality.
@@ 1 Introduction
Score-based generative models (Song et al., 2021), also termed diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), have shown remarkable potential for the generation of images (Dhariwal & Nichol, 2021; Rombach et al., 2022), videos (Ho et al., 2022), text (Li et al., 2022; Dieleman et al., 2022; Wu et al., 2023), molecules (Hoogeboom et al., 2022), and many other highly complex data structures with continuous features. The framework has since been adapted to categorical data in various ways, including discrete diffusion processes (Austin et al., 2021; Hoogeboom et al., 2021), diffusion in continuous embedding space (Dieleman et al., 2022; Li et al., 2022; Regol & Coates, 2023; Strudel et al., 2022), and others (Campbell et al., 2022; Meng et al., 2022; Sun et al., 2023). Diffusion models which include both, continuous and categorical features alike, build directly on advances from the image domain (Kim et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Jolicoeur-Martineau et al., 2024) and thus, are not designed to deal with challenges specific to mixed-type data: The different diffusion processes and their losses are neither aligned nor balanced across data types, and do not scale to larger datasets and/or features with a greater number of categories. Models that naively combine different losses to integrate distinct generative processes may suffer from implicitly favoring the sample quality of some features or data types over others (Ma et al., 2020).
@@
A crucial component in score-based generative models is the noise schedule (Kingma et al., 2022; Chen et al., 2022; Chen, 2023; Jabri et al., 2022; Wu et al., 2023). Typical noise schedules for image and text data are designed to focus model capacity on the noise levels most important to sample quality (Nichol & Dhariwal, 2021; Karras et al., 2022), while others attempt to learn the optimal noise schedule (Dieleman et al., 2022; Kingma et al., 2022). For mixed-type tabular data, existing approaches often combine distinct diffusion processes for the continuous and discrete features to derive a joint model (Kotelnikov et al., 2023; Lee et al., 2023). However, noise schedules are not directly transferable from one data modality to another and therefore, using specifications from image or text domain models is not optimal: First, the inherently different diffusion processes make it difficult to balance the noise schedules across features and feature types, and negatively affect the allocation of model capacity across timesteps. Further, and most importantly, the domain, nature and marginal distribution can vary significantly across features (Xu et al., 2019). For instance, any two continuous features may be subject to different levels of discretization or different bounds, even after applying common data pre-processing techniques; and any two categorical features may differ in the number of categories, or the degree of imbalance. The high heterogeneity and lack of balancing warrants a rethinking of fundamental parts of the diffusion framework, including the noise schedule and the effective combination of diffusion processes for different data types.
@@
In this paper, we introduce Continuous Diffusion for mixed-type Tabular Data (CDTD) to address the aforementioned shortcomings. We combine score matching (Hyvärinen, 2005) with score interpolation (Dieleman et al., 2022) to derive a score-based model that pushes the diffusion process for categorical data into embedding space, and uses a Gaussian diffusion process for both continuous and categorical features. This way, the different noise processes become directly comparable, easier to balance, and enable the application of, for instance, classifier-free guidance (Ho & Salimans, 2022), accelerated sampling (Lu et al., 2022), and other advances, to mixed-type tabular data.
@@
We counteract the high feature heterogeneity inherent to data of mixed-type with distinct feature or type-specific adaptive noise schedules. The learnable noise schedules allow the model to directly take feature or type heterogeneity into account during both training and generation, and thus avoid the reliance on image or text-specific noise schedule designs. Moreover, we propose a diffusion-specific loss normalization and initialization scheme to homogenize different data types and their losses effectively. Our improvements ensure a better allocation of the model’s capacity across features, feature types and timesteps, and yield high quality samples of tabular data. CDTD outperforms state-of-the-art baseline models across a diverse set of sample quality metrics as well as computation time for data sets with an arbitrary number of categories and data points. Our experiments show that CDTD captures feature correlations exceptionally well, and that explicitly allowing for data-type heterogeneity in the noise schedules benefits sample quality.
@@
In sum, we make several contributions specific to diffusion probabilistic modeling of tabular data:
@@
  • We propose a joint continuous diffusion model for both continuous and categorical features such that all noise distributions are Gaussian.
@@
  • We balance model capacity across continuous and categorical features with a novel and effective loss calibration, an adjusted score model initialization and type or feature-specific noise schedules.
@@
  • We extend the idea of timewarping and propose a functional form to efficiently learn adaptive noise schedules, and to allow for exact evaluation and easy incorporation of prior information on the relative importance of noise levels.
@@
  • We drastically improve the scalability of tabular data diffusion models to features with a high number of categories.
  • We boost the quality of the generated samples with adaptive, feature or type-specific noise schedules.
@@
  • Our CDTD model allows the first-ever use of advanced techniques, like classifier-free guidance, for mixed-type tabular data directly in data space.
@@ 2 Score-based Generative Framework
We start with a brief outline of the score-based frameworks for continuous and categorical features. Next, we combine these into a single diffusion model to learn the joint distribution of mixed-type data.
@@ 2.1 Continuous Features
In this paper, we use the EDM formulation (Karras et al., 2022), that is, ${\mathbf{f}}(\cdot,t)={\bm{0}}$ and $g(t)=\sqrt{2[\frac{d}{dt}\sigma(t)]\sigma(t)}$ such that $p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{x}}_{0},\sigma^{2}(t)I_{K_{\text{cont}}})$. We start the reverse process with sampling ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{K_{\text{cont}}})$ for $\sigma^{2}(1)$ being sufficiently large and $\mathbb{E}[{\mathbf{x}}_{0}]={\bm{0}}$. We then gradually guide ${\mathbf{x}}_{1}$ towards high density regions in the data space with ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},t)$ replacing the unknown, true score function in Equation 2. In practice, ODE or predictor-corrector samplers can be used for this iterative denoising process (Song et al., 2021).
@@ 2.2 Categorical Features
For categorical features, the score function is undefined and the score-based generative framework cannot be directly applied. Therefore, we use score interpolation (Dieleman et al., 2022) and push the diffusion of categorical data into Euclidean embedding space. This allows us to impose a Gaussian noise distribution on categorical and continuous features.
Let $x_{\text{cat}}^{(j)}\in\{1,\dots,C_{j}\}$ be the $j$-th categorical feature with $C_{j}$ possible classes. The $i$-th class is associated to a distinct, trainable $d$-dimensional embedding vector ${\mathbf{e}}_{i}^{(j)}\in{\mathbb{R}}^{d}$. The embedding vector corresponding to the ground truth class is ${\mathbf{x}}_{0}^{(j)}\in\{{\mathbf{e}}_{1}^{(j)},\dots,{\mathbf{e}}_{C_{j}}^{(j)}\}$. We omit the superscript $j$ in what follows. The noisy variant at time $t$ is ${\mathbf{x}}_{t}\sim p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{x}}_{0},\sigma^{2}(t)I_{d})$ similar to before. Given ${\mathbf{x}}_{t}$ and $t$, $\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})]$ minimizes Equation 3, and we obtain
@@
$\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\frac{1}{\sigma^{2}(t)}[\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]-{\mathbf{x}}_{t}]\ .$ (4)
@@
We compute $\hat{{\mathbf{x}}}_{0}=\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]$ as an average over the $C_{j}$ possible embedding vectors weighted by their probability. This yields a score function estimate for a single categorical feature. Since $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)=p(x_{\text{cat}}^{(j)}=i|{\mathbf{x}}_{t},t)$, an estimate of $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)$ is obtained via a classifier that predicts $C_{j}$ class probabilities for the $j$-th feature and is trained via cross-entropy (CE) loss.
We start generating a categorical feature from an embedding vector ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{d})$. We then use score interpolation (Equation 4) to iteratively denoise this embedding vector. After the final reverse step, i.e., when  $t\approx 0$, we can directly infer the generated class from the feature-specific class probabilities, $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)$.
@@
This framework can easily be extended to multiple categorical features. All embeddings are trained alongside the model. The model fully takes feature-specific uncertainty at intermediate timesteps into account (as for each feature it only has to commit to a category at the final step of generation) which improves the consistency of the generated samples (Dieleman et al., 2022). Therefore, CDTD allows to capture subtle dependencies both within and across data types more accurately.
@@ 3 Method
In short, we combine score matching (Equation 3) with score interpolation (Equation 4) to model the joint distribution of mixed-type data. Next, we discuss the important components of our method. In particular, the combination of the different losses for score matching and score interpolation, initialization and loss weighting concerns, and the adaptive type- or feature-specific noise schedule designs.
@@ 3.1 General Framework
Figure 1 gives an overview of our Continuous Diffusion for mixed-type Tabular Data (CDTD) framework. The score model is conditioned on (1) all noisy continuous features, (2) the noisy embeddings of all categorical features in Euclidean space, and (3) the timestep $t$ which reflects potentially feature-specific, adaptive noise levels $\sigma_{\text{cont},i}$ and $\sigma_{\text{cat},j}$ for all $i$ and $j$. Additional conditioning information, such as the target feature for classification tasks, are straightforward to add. Note that while the Gaussian noise process acts directly on the continuous features, it acts on the embeddings of the categorical features. This way, we ensure a common continuous noise process for both data types.
During training, the model predicts the ground-truth value for continuous features and the class-specific probabilities for categorical features. During generation, we concatenate the score estimates, $\hat{s}_{\text{cont}}^{(i)}$ and $\hat{s}_{\text{cat}}^{(j)}$, for all features $i$ and $j$, and pass them to an ODE solver together with $\sigma_{\text{cont},i}$ and $\sigma_{\text{cat},j}$, the noise levels retrieved by transforming linearly spaced timesteps with the learned adaptive noise schedules. Further details on the implementation and sampling are provided in Appendix J and Appendix K, respectively.

### v4

@@ Abstract
Score-based generative models, commonly referred to as diffusion models, have proven to be successful at generating text and image data. However, their adaptation to mixed-type tabular data remains underexplored. In this work, we propose CDTD, a Continuous Diffusion model for mixed-type Tabular Data. CDTD is based on a novel combination of score matching and score interpolation to enforce a unified continuous noise distribution for both continuous and categorical features. We explicitly acknowledge the necessity of homogenizing distinct data types by relying on model-specific loss calibration and initialization schemes. To further address the high heterogeneity in mixed-type tabular data, we introduce adaptive feature- or type-specific noise schedules. These ensure balanced generative performance across features and optimize the allocation of model capacity across features and diffusion time. Our experimental results show that CDTD consistently outperforms state-of-the-art benchmark models, captures feature correlations exceptionally well, and that heterogeneity in the noise schedule design boosts sample quality. Replication code is available at https://github.com/muellermarkus/cdtd.
@@ 1 Introduction
Score-based generative models (Song et al., 2021), also termed diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), have shown remarkable potential for the generation of images (Dhariwal & Nichol, 2021; Rombach et al., 2022), videos (Ho et al., 2022), text (Li et al., 2022; Dieleman et al., 2022; Wu et al., 2023), molecules (Hoogeboom et al., 2022), and other highly complex data structures with continuous features. The framework has since been adapted to categorical data in various ways, including discrete diffusion processes (Austin et al., 2021; Hoogeboom et al., 2021), diffusion in continuous embedding space (Dieleman et al., 2022; Li et al., 2022; Regol & Coates, 2023; Strudel et al., 2022), and others (Campbell et al., 2022; Meng et al., 2022; Sun et al., 2023). However, their adaptation to mixed-type tabular data, which includes both continuous and categorical features, remains under explored. Existing models build directly on advances from the image domain (Kim et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Jolicoeur-Martineau et al., 2024) and, therefore, are not designed to deal with challenges specific to mixed-type tabular data: The type-specific diffusion processes and their losses are neither aligned nor balanced. A naive combination of different losses can cause the generative model to favor the sample quality of some features or data types over others (Ma et al., 2020). Furthermore, tabular data often includes categorical features of high cardinality. However, existing diffusion models for tabular data typically rely on a discrete diffusion framework to model one-hot-encoded categorical features (e.g., Kotelnikov et al., 2023; Lee et al., 2023). As a consequence, these models do not scale well and fail to capture the full uncertainty during the denoising process, as a data sample can never be ‘in-between’ categories.
@@
Noise schedules determine the amount of noise at each diffusion timestep and are typically defined manually (Nichol & Dhariwal, 2021; Karras et al., 2022), but can also be learned (Dieleman et al., 2022; Kingma et al., 2021). They are a crucial component in score-based generative models (Kingma et al., 2021; Chen et al., 2023; Chen, 2023; Jabri et al., 2023; Wu et al., 2023) as they aim to focus the model capacity on the noise levels most important to sample quality. However, despite their importance, previous work on diffusion models for tabular data simply adopted noise schedules from the image domain, which is not optimal. First, diffusion models for mixed-type tabular data inherently rely on combining type-specific diffusion processes. This makes it important, but difficult, to balance noise schedules across feature types. Unbalanced noise schedules negatively affect the allocation of model capacity. For instance, both TabDDPM (Kotelnikov et al., 2023) and CoDi (Lee et al., 2023) use the discrete multinomial diffusion framework (Hoogeboom et al., 2021) to model categorical features. This induces different types of noise for continuous and categorical features, making alignment or comparison of noise schedules impossible. Second, the domain, nature, and marginal distribution can vary significantly across features (Xu et al., 2019). For instance, any two continuous features may be subject to different levels of discretization or skewness, even after applying common data preprocessing techniques; and any two categorical features may differ in the number of categories, or the degree of imbalance. The high heterogeneity and lack of balance warrant a rethinking of fundamental parts of the diffusion framework, including the noise schedule and the effective combination of diffusion processes for different data types, rather than simply copying what has been working for images.
@@
In this paper, we introduce Continuous Diffusion for mixed-type Tabular Data (CDTD) to address the aforementioned shortcomings. We combine score matching (Hyvärinen, 2005) with score interpolation (Dieleman et al., 2022) to derive a score-based model that pushes the diffusion process for categorical data into embedding space, and thus enables a Gaussian diffusion process for both continuous and categorical features. This way, the different noise processes become directly comparable, easier to balance, and enable the application of, for instance, classifier-free guidance (Ho & Salimans, 2022), accelerated sampling (Lu et al., 2022), and other advances to mixed-type tabular data.
@@
We counteract the high feature heterogeneity inherent to mixed-type tabular data with distinct feature- or type-specific adaptive noise schedules. The learnable noise schedules allow the model to directly take feature or type heterogeneity into account during both training and generation, and thus avoid the reliance on image-specific noise schedule designs. Likewise, Shi et al. (2024) also propose the use of feature-specific noise schedules for tabular data but do not account for the necessity of homogenizing different data types. In contrast, we propose a diffusion-specific loss calibration and initialization scheme for effective data type homogenization. These improvements ensure a better allocation of our model’s capacity across features, feature types and timesteps, and yield high quality samples. CDTD outperforms state-of-the-art baseline models across a diverse set of sample quality metrics and datasets as well as computation time. Our experiments show that CDTD captures feature correlations exceptionally well, and that explicitly allowing for data-type heterogeneity in the noise schedules benefits sample quality.
@@
In sum, we make several contributions specific to score-based modeling of tabular data:
@@
  • We propose a unified continuous diffusion model for both continuous and categorical features such that all noise distributions are Gaussian.
@@
  • We balance model capacity across continuous and categorical features with a novel loss calibration, an adjusted score model initialization and adaptive type- or feature-specific noise schedules.
@@
  • We suggest a novel functional form to efficiently learn adaptive noise schedules, and to allow for an exact evaluation and incorporation of prior information on the relative importance of noise levels.
@@
  • We drastically improve the scalability of tabular data diffusion models to high-cardinality features.
@@
  • We are the first to enable the use of advanced techniques, like classifier-free guidance, for mixed-type tabular data directly in data space as opposed to a latent space.
@@ 2 Score-based Generative Framework
We start with outlining the score-based frameworks for continuous and categorical features. Next, we combine these into a single, unified model to learn the joint distribution of mixed-type tabular data.
@@ 2.1 Continuous Features
In this paper, we use the EDM formulation (Karras et al., 2022), that is, ${\mathbf{f}}(\cdot,t)={\bm{0}}$ and $g(t)=\sqrt{2[\frac{\mathrm{d}}{\mathrm{d}t}\sigma(t)]\sigma(t)}$ such that $p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{x}}_{0},\sigma^{2}(t){\mathbf{I}}_{K_{\text{cont}}})$. We standardize ${\mathbf{x}}_{0}$ to zero mean and unit variance. Then, for a sufficiently large $\sigma^{2}(1)$, we can start the reverse process with sampling ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1){\mathbf{I}}_{K_{\text{cont}}})$. We then gradually guide ${\mathbf{x}}_{1}$ towards high density regions in the data space with ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},t)$ replacing the unknown, true score function in Equation 2.
@@ 2.2 Categorical Features
Let $x_{\text{cat}}^{(j)}\in\{1,\dots,C_{j}\}$ be the $j$-th categorical feature with $C_{j}$ distinct classes. We learn a feature-specific encoder to represent each category $c$ as a $d$-dimensional vector ${\mathbf{e}}_{x_{\text{cat}}}^{(j)}=\text{Enc}_{j}(x_{\text{cat}}^{(j)})$. Further, let ${\mathbf{x}}_{0}^{(j)}\in\{{\mathbf{e}}_{1}^{(j)},\dots,{\mathbf{e}}_{C_{j}}^{(j)}\}$ be the noiseless embedding at $t=0$. To unify the diffusion frameworks for categorical and continuous data as much as possible, we base both on the same Gaussian-type noise. Thus, instead of adding noise to the categorical variable directly, we add noise to the embedding ${\mathbf{x}}^{(j)}_{t}\sim p_{0t}({\mathbf{x}}_{t}^{(j)}|{\mathbf{x}}^{(j)}_{0})=\mathcal{N}({\mathbf{x}}^{(j)}_{t}|{\mathbf{x}}^{(j)}_{0},\sigma^{2}(t)I_{d})$ such that ${\mathbf{x}}^{(j)}_{1}\sim p_{1}({\mathbf{x}}^{(j)})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{d})$, analogous to score matching.
For categorical data, denoising score matching (Equation 3) is not directly applicable to learn $\nabla_{{\mathbf{x}}^{(j)}_{t}}\log p_{0t}({\mathbf{x}}^{(j)}_{t}|{\mathbf{x}}^{(j)}_{0})$, since the score can only take on $C_{j}$ distinct values. To proceed, we transform the score matching approach into a discrete choice problem. Note that for a given $t$ and ${\mathbf{x}}^{(j)}_{t}$ it is sufficient to find $\mathbb{E}_{\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)}[\nabla_{{\mathbf{x}}^{(j)}_{t}}\log p_{0t}({\mathbf{x}}^{(j)}_{t}|{\mathbf{x}}^{(j)}_{0})]$ as it minimizes Equation 3. Presuming Gaussian noise, we have
@@
$\mathbb{E}_{\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)}\Bigl{[}\nabla_{{\mathbf{x}}^{(j)}_{t}}\log p_{0t}({\mathbf{x}}^{(j)}_{t}|{\mathbf{x}}^{(j)}_{0})\Bigr{]}=\frac{1}{\sigma^{2}(t)}\Bigl{[}\mathbb{E}_{\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)}[{\mathbf{x}}^{(j)}_{0}]-{\mathbf{x}}^{(j)}_{t}\Bigr{]}\ .$ (4)
@@
We can thus approximate the score by computing the probability-weighted average of the $C_{j}$ possible embedding vectors, that is, $\hat{{\mathbf{x}}}^{(j)}_{0}=\mathbb{E}_{\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)}[{\mathbf{x}}^{(j)}_{0}]$. Since $\mathrm{P}({\mathbf{x}}^{(j)}_{0}={\mathbf{e}}^{(j)}_{x_{\text{cat}}}|{\mathbf{x}}^{(j)}_{t},t)=\mathrm{P}(x_{\text{cat}}^{(j)}=c|{\mathbf{x}}^{(j)}_{t},t)$, we can estimate $\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)$ via a classifier that predicts the $C_{j}$ class probabilities and is trained to minimize the cross-entropy (CE). This procedure effectively interpolates between the $C_{j}$ ground-truth embeddings ${\mathbf{x}}_{0}^{(j)}$ and is therefore known as score interpolation (Dieleman et al., 2022).
@@
This framework easily extends to multiple features. Most importantly, $\text{Enc}_{j}$ is trained alongside the model such that ${\mathbf{x}}_{0}^{(j)}$ is directly optimized for denoising the data. During sampling, the model only has to commit to a category at the final step of generation, i.e., we allow for a smooth, continuous transition between states. This is unlike multinomial diffusion (Hoogeboom et al., 2021), which imposes discrete transitioning steps and is the framework used by several existing diffusion models for tabular data (Kotelnikov et al., 2023; Lee et al., 2023). By defining diffusion for categorical data in embedding space, we allow our model to take uncertainty at intermediate timesteps fully into account, which improves the consistency of the generated samples (Dieleman et al., 2022). Thus, CDTD can more accurately capture subtile dependencies both within and across data types.
@@ 3 Method
To model the joint distribution of mixed-type data, we combine score matching (Equation 3) with score interpolation (Equation 4). Next, we discuss the important components of our method. In particular, the combination of the different losses for score matching and score interpolation, initialization and loss weighting concerns, and the adaptive type- or feature-specific noise schedule designs.
@@ 3.1 General Framework
Figure 1 gives an overview of our Continuous Diffusion for mixed-type Tabular Data (CDTD) framework. The score model is conditioned on (1) the noisy continuous features, (2) the noisy embeddings of categorical features, and (3) the timestep $t$. It predicts the ground-truth value $x_{\text{cont}}^{(i)}$ for continuous features and the class-specific probabilities $\mathrm{P}(x_{\text{cat}}^{(j)}=c)$ for categorical features. Additional conditioning information is straightforward to add. Note that while the Gaussian noise process acts directly on continuous features, it acts on the embedded categorical features ${\mathbf{x}}_{0}^{(j)}$. This way, we ensure a unified continuous noise process for both data types. Further details on the implementation are provided in Appendix K. During generation, we concatenate the score estimates, $\hat{s}_{\text{cont}}^{(i)}$ and $\hat{s}_{\text{cat}}^{(j)}$, for all features $i$ and $j$ before passing them to an ODE solver. Our model is the first to utilize feature- or type-specific noise schedules also during the generation of tabular data, which allows the sampler to take feature- or type-specific steps. Details on the sampling process and the algorithm are given in Appendix L.